OpenStudy (anonymous):

Coach Johnson is 10 years older than Tim. When Tim started football 3 years ago, Coach Johnson was twice his age. How old is Tim now?

OpenStudy (anonymous):

13= Tim's age now, so coach is 23; three years ago they were 10 & 20

OpenStudy (anonymous):

solve the system: johnson - J tim - T J=T+10 J-3=2(T-3)